# Did I Vet You Before? Assessing the Chrome Web Store Vetting Process through Browser Extension Similarity

### v1

@@ 1. Introduction
Over the past decade, web browsers have become an indispensable part of both the desktop and mobile ecosystems. As more and more traditional applications have migrated to online services, users spend more time in their browsers, increasing the need for extended and more customization features in an essential piece of software that they use every day. Chromium-based browsers, most notably Google Chrome, allow this customization through browser extensions: third-party programs built with web technologies that can interact with the browser to modify the user experience. Some popular types of extensions include ad blockers, grammar checkers, password managers, and, more recently, AI assistants.
@@
The undisputed marketplace for browser extensions for Chromium browsers is the Chrome Web Store (CWS), Google’s official marketplace launched back in 2011 (Kay and Boodman, 2011). The CWS plays a vital role in the browser ecosystem as many users inherently trust the extensions included on its catalogue, but evidence shows that malicious and deceptive ones eventually get published (Palant, 2023; Titterington, 2023). To mitigate the risk of harmful extensions being distributed through the official store, publishers must adhere to the CWS Program Policies, which govern what content can be uploaded. If an infringing extension is published, Google can take it down at its discretion during the vetting process and, in some cases, remove it from the browsers that installed it.
Research Questions and Key Findings. Given the ever-increasing importance of web browsers—and, indirectly, the CWS as the largest distribution platform—, it is important to study the prevalence of infringing content in the CWS and how effective the CWS vetting process is at dealing with it. In the past, the CWS has received reports from security vendors and researchers notifying them of infringing content (Palant, 2023; Titterington, 2023; ReasonLabs Research Team, 2023). There is also evidence of Google themselves proactively vetting extensions that do not belong in the store (Jagpal et al., 2015). This leads us to our first research question:
@@
RQ1. How effective is the CWS vetting process at detecting and rapidly removing infringing extensions?
To answer this question we follow the data collection and analysis methodology shown in Figure 1. We compile a dataset of 366,617 extensions downloaded from the CWS during 4 years. We enrich this dataset with vetting labels assigned by Google moderators that indicate which extensions have been vetted and for what reason. We then design and develop SimExt, a novel methodology for finding similarly behaving extensions that leverages $(i)$ static and dynamic analysis for extracting syntactic and behavioral features, and $(ii)$ Natural Language Processing (NLP) and vector embeddings for computing similarity. Using SimExt, we perform cluster analysis to find unvetted extensions that are very similar to vetted ones, as these are natural candidates for being infringing items that were overlooked by the vetting process. We find that 17% of the infringing extensions flagged by our pipeline have not been removed from the CWS. Interestingly, 86% of infringing items are republished extensions, i.e., extensions that are either identical or extremely similar to items that were previously vetted, suggesting that the CWS vetting process does not adequately learn from experience. We also find that 11% of the infringing extensions that remain published at the end of crawl come from repeat offenders (publishers with known vetted items), suggesting that the CWS does not suspend these accounts according to their own policies. Our analysis also reveals that the process is excessively slow in vetting extensions, taking months or even years from the publication of an infringing version to its removal from the store.
Characterizing the trends and challenges that the CWS faces can help improve the effectiveness of the vetting process. To do so, we need to understand what types of infringing content are most commonly uploaded to the store, especially those that escape the scrutiny of moderators, so that they can focus on identifying and taking them down. This challenge is the motivation for the second research question studied in this work:
RQ2. What are the main kinds and the key features of infringing extensions uploaded to the CWS?
We address this problem by analyzing the top infringing clusters with the largest number of extensions and manually assigning them a type or kind of content. We find that 83% of the extensions in these clusters are New Tab Extensions (NTEs), which are items that override the webpage that loads when a new tab is opened. The cluster analysis also reveals the presence of spam and no-content extensions, both vetted and still published in the CWS. In order to better understand the labeling process, we explore the relationship between the CWS vetting labels and the behavior of an extension. Our analysis finds that labels are somewhat inconsistent, especially in the case of NTEs.
An important class of infringing extensions are those labeled by CWS analysts as malware. Previous research has studied browser extension malware, focusing primarily on how to detect it (Kapravelos et al., 2014; Jagpal et al., 2015; Aggarwal et al., 2018; Pantelaios et al., 2020; Picazo-Sanchez et al., 2022). One question that remains unexplored is understanding the types of malicious extensions that target the CWS, including their behavior, top families, impact on users, and challenges for detection. This constitutes our third research question:
@@
RQ3. What malware families target the CWS and what threat intelligence is available about them?
@@
To answer this question, we scan the 5,647 extensions labeled as malware by the CWS in VirusTotal and analyze the results. Remarkably, 95% of all samples were not previously seen by VirusTotal and 3% where known but have zero detections. This finding suggests that the majority of malicious extensions detected in the CWS are unknown to the threat intelligence community, and that for others there is a significant discrepancy between the criteria of CWS moderators and the detection capabilities of malware detection engines. The analysis of malicious extensions for which we obtain detection reports reveals that the malware lineage analysis conducted by security vendors is very poor, with the malware family label missing in most cases and being very general in others.
@@
Disclosure. We reported to Google a subset of manually reviewed, infringing extensions found by our pipeline. Most importantly, these include a cluster comprised of items published from corporate Google accounts. While most of the latter have been removed, at the time of this writing we have not received a response.
@@
Research Artifacts. We provide several artifacts for reproducibility at https://zenodo.org/records/10977708. These artifacts include the compiled list of extension pairs used as ground truth for evaluation, as well as metadata and high-dimensional embeddings for all the infringing extensions found by our tool.
2. Background
Figure 1. Data collection and analysis methodology pipeline.
@@ 2.1. Browser Extensions
Browser extensions are third-party programs that enhance or modify the functionality of the browser, with ad blockers being one of their most recognizable and widely-used examples. Extensions are written using a combination of HTML, CSS and JavaScript, and may contain other resources commonly found in web applications, such as images or fonts (MDN contributors, 2024a). Extensions are packaged as CRX files, bundling all the source code and assets needed to run them into a single file, thus easing distribution. CRX packages are regular ZIP archives with an additional header that ensures the integrity and authenticity of its contents (Pawlicki, 2017).
@@
As in the case of other popular platforms like Android, all extensions have a mandatory manifest.json file that dictates their functionality and behavior. This manifest is a JSON document that specifies basic metadata about the extension, like its name and version. Optionally, it may also declare permissions, service workers (which run code in the background), content scripts (which run code in open tabs), and overridden pages (e.g., to modify the default URL when opening a new tab), among other elements (Google Chrome Developers, 2012a). Two popular types of extensions are themes and New Tab Extensions (NTEs). A theme is a type of extension that changes the way the browser looks and does not contain any source code, i.e., all theme configuration is in the manifest itself. An NTE is an extension that changes the default webpage that will load every time a new tab is opened.
@@
Extensions communicate with the browser using Web APIs (MDN contributors, 2024b) and Extension APIs (Google Chrome Developers, 2024a). These APIs provide extensions with a broad range of powerful capabilities, allowing them to access and manipulate the Document Object Model (DOM) of webpages, and to access geolocation data, browsing history, and other sensitive information stored in the browser. Some of these APIs are restricted by default for security reasons, and can only be accessed by declaring the appropriate permissions in the manifest.
2.2. CWS Program Policies
@@
CRX packages can be published on the Chrome Web Store (CWS), Google’s official distribution platform for browser extensions that run on Google Chrome (Google Chrome Developers, 2024b). Users of other Chromium-based browsers, such as Microsoft Edge, Brave, or Opera, can also install extensions from this source as they run the same compatible engine. In fact, Brave even recommends that users download extensions from this platform in the absence of their own official store (Brave Software, Inc., 2023). As such, policing the content that is uploaded to the CWS is of great importance because of the impact that harmful extensions can have given the large audience of the store.
@@
The CWS requires all developers to comply with a set of policies that include security and privacy guidelines. These policies prohibit the publication of extensions with a malicious purpose and extensions that facilitate unauthorized access to restricted or copyrighted content (Google Chrome Developers, 2022b). They also ask developers not to obfuscate code or conceal functionality, to request the minimum set of permissions needed, and forbid submitting multiple extensions that provide duplicate or very little functionality (Google Chrome Developers, 2022a, 2020c, c). According to Google’s policies, repeated violations of these policies will result in the suspension of the publisher account (Google Chrome Developers, 2022d).
3. Data Collection
@@
We implemented a purpose-specific crawler to download the resources shown in Figure 1, namely CRX packages and their associated metadata, from Google’s Chrome Web Store (Google Chrome Developers, 2024b). Our crawling strategy consists of three steps. First, we prepare an initial seed of extensions from the store’s sitemap, the reels of popular items from each category, and the featured extensions from the home page. We then visit each extension page, extract metadata (e.g., name, description, number of installs, last modification date) and download its associated assets (i.e., CRX package, icon and tile). Finally, we look for non-visited extensions appearing as “related” or “recommended” in the store page of each visited extension and add them to the queue of items to crawl. Previous efforts demonstrate the effectiveness of this approach (Fass et al., 2021; Laperdrix et al., 2021), which takes inspiration from strategies successfully used to crawl the Google Play Store (Wang et al., 2018; Petsas et al., 2013).
@@
We run the crawler daily to detect new and removed extensions, and keep track of changes being pushed to the store such as publishers uploading an updated version of an existing extension. For simplicity, we set the visitor’s country code to the United States when crawling the store. Although developers can opt-out of publishing extensions in certain countries or even entire regions (Google Chrome Developers, 2020b), this decision should not limit our coverage as we can still download any non-private CRX package regardless, and because the used sitemap contains publicly listed extensions worldwide. We only run our data collection pipeline once a day and rely on the HTTP caching headers provided by CWS to responsibly limit our impact on Google’s servers. We note that in October 2023, we transitioned to crawling the new version of the CWS, coinciding with its first public release (Ismail, 2023). We made this decision in preparation for the announced retirement of the previous version in January 2024. While we experienced no issues during this migration, we also kept crawling the legacy store for a few more weeks in the event something went wrong.
@@
In addition to crawling CRX packages and store metadata, we also fetch vetting labels issued by Google from a separate endpoint. These labels indicate whether an extension was taken down from the CWS instead of willingly being unpublished by its developer, as well as the reason behind the takedown (e.g., flagged as malware), if any. Vetting labels are used by Google Chrome since version 117 to protect users of vetted extensions by proactively disabling malware and displaying warnings in the browser extensions page (Dunk, 2023). As far as we can tell, this is the first study to crawl and analyze this data.
3.1. Dataset
@@
We crawled the CWS for 4 years from March 2020 to March 2024. We compiled a historical dataset of 366,617 extensions, containing 902,532 different updates (or extension versions) that were available for download at some point in the store. Some of the extensions in our datasets were published as early as December 2009. Of all these versions, we lack the CRX package for 28,405 paid extensions that did not change their pricing model after payments were deprecated in the CWS in mid-2020 (Google Chrome Developers, 2020a). Our dataset also contains 53,132 browser themes, which are distributed in the same form as extensions but do not have the ability to execute code. To the best of our knowledge, this is the most comprehensive dataset of browser extensions to-date, with over 156k more extensions than the previous large-scale work that crawled the CWS (Pantelaios et al., 2020).
@@
Figure 2. Daily volume of published extensions (solid, left axis) and accumulated dataset size (dashed, right axis).
@@
Vetting Labels. Using the vetting labels that we crawl for removed extensions, we can infer whether they were unpublished by the developer or taken down by Google. Vetted extensions have a type of either Malware, Policy Violation, or Minor Policy Violation, whereas items that were unpublished at the request of their developers have a type of None. Figure 3 shows the ratio of extension labels found in our dataset at a given point in time since we first started crawling. As of March 2024, 52% of all the extensions we collected have been removed from the store. More interestingly, 79% of the removed items are infringing extensions that were taken down. This preliminary characterization of the dataset motivates us to delve deeper into the dynamics of extension removals and the vetting process of the CWS.
Figure 3. Daily ratio of vetting labels provided by CWS.
4. SimExt: A Methodology for Finding Similar Extensions
While developing a fully automated vetting process is very challenging, having an automatic approach to detect similar extensions at scale is valuable to assist analysts, as it can be used to find items that resemble previously vetted ones or clones to currently published content. Our key working hypothesis is that if an extension is taken down from the store because it violates a program policy, then other extensions that are highly similar are candidates to be vetted as well. Even in those cases where the behavior of an extension is not necessary harmful, finding clones is still valuable given the CWS ban on repetitive content (Google Chrome Developers, 2020c). We acknowledge that legitimate extensions cloned by malicious actors are an exception to this rule. For example, the vetting of malicious extensions reusing MetaMask’s codebase should not result in MetaMask being taken down. We consider that these cases should be manually filtered out, and that a tool that provides detailed similarity reports between two extensions can be extremely helpful to assist the store vetting process. Section 6 provides examples of published infringing extensions found by our tool to support this claim.
Since there is currently no tool that identifies similar extensions for a given input item, we designed and built SimExt, our own state-of-the-art solution for querying browser extensions based on similarity. We design it as a multi-step pipeline to automatically extract meaningful behavioral features from all extensions in our dataset and then group them into clusters based on their similarity. This approach is different from previous work that classifies extensions as either benign or malicious (Jagpal et al., 2015; Aggarwal et al., 2018; Pantelaios et al., 2020). Moreover, to the best of our knowledge, our work is the first one to leverage Natural Language Processing (NLP) and vector embeddings to effectively and accurately find similar extensions at scale. Figure 1 provides a summary of the pipeline steps described below.
4.1. Feature Extraction

### v2

@@ I Introduction
Over the past decade, web browsers have become an indispensable tool for both desktop and mobile users. As traditional applications migrate to web-based services, users spend more time in their browsers, increasing the need for extended and more customization features. Chromium-based browsers—notably, Google Chrome—allow this customization through browser extensions: third-party programs built with web technologies that interact with the browser to add or modify features, and enhance the user experience. Some popular types of extensions include ad blockers, grammar checkers, password managers, and, more recently, AI assistants.
@@
The undisputed marketplace for browser extensions for Chromium browsers is the Chrome Web Store (CWS), Google’s official marketplace which launched back in 2011 [30]. The CWS plays a vital role in the browser ecosystem as many users inherently trust the extensions included on its catalog. Similar to the case of the Google Play Store for Android apps, the CWS implements vetting processes to detect and mitigate the risk of harmful and deceptive extensions being distributed through the official store. Extension publishers must adhere to the CWS Program Policies, which govern what content can be uploaded and distributed [23]. These policies forbid not only the publication of harmful content, but also mandate that extensions must disclose their behavior, implement a minimum of functionality, and not duplicate existing content, among other requirements. If an infringing extension (i.e., one that does not abide by the CWS policies) is published, Google may delist it at its discretion during the vetting process and, in some cases, even remove it from browsers that have it installed.
In the past, the CWS has been under scrutiny from security experts and researchers. Various reports and research studies identified infringing content being distributed on the CWS, focusing mostly on potentially harmful extensions [46, 61, 53] and in methods to detect their presence [29, 28, 1, 48, 51]. While there is evidence of Google proactively vetting extensions from its official store [28], the effectiveness of the vetting process at detecting infringing content remains unexplored. Recent research has focused on characterizing the content of the store and its trends [26], including the presence of extensions with vulnerabilities or a similar code base. However, a knowledge gap remains when it comes to understanding Google’s policing of the platform. Characterizing the challenges that the CWS vetting process faces in practice can help identify its pitfalls and improve its effectiveness. To fill this gap, this paper aims to answer the following research questions:
  • RQ1.
@@
    How effective is the CWS vetting process at detecting and rapidly removing infringing extensions?
  • RQ2.
    What are the main kinds and the key features of infringing extensions uploaded to the CWS?
  • RQ3.
@@
    What malware families target the CWS and what threat intelligence is available about them?
To answer these questions, we develop the scalable data collection and hybrid analysis pipeline for browser extensions shown in Figure 1. We compile a dataset of 366,617 extensions downloaded from the CWS during 4 years. We enrich this dataset with vetting labels assigned by Google moderators to removed CWS items. These labels indicate which extensions have been taken down from the store and for what reason, thus providing ground truth about the vetting process. To the best of our knowledge, this is the first work that studies the CWS vetting process using ground truth about removed content. We then design and develop SimExt, a novel methodology for finding similarly behaving extensions that leverages $(i)$ static and dynamic analysis for extracting syntactic and behavioral features; and $(ii)$ Natural Language Processing (NLP) and vector embeddings for computing similarity. To address the existing technical gap in feature extraction from browser extensions, we develop and release Fakeium [42], an open source dynamic analysis sandbox that follows an implementation inspired by concolic execution to elicit behaviors and overcome the limitations of static analysis and fuzzy-hash-based methods [26]. Fakeium is very scalable and can process an extension in mere seconds, which makes it suitable for large-scale measurements such as the one conducted in this work. To generate the vector embeddings, we use a novel approach based on Zero-Short Learning (ZSL) which, unlike previous work, does not require retraining the model when new malicious extensions are uploaded to the store [28, 1, 48, 51].
Using SimExt, we perform a large-scale behavioral cluster analysis to find published extensions that are very similar to vetted ones, as these are natural candidates for being infringing items that were overlooked by the vetting process. Equipped with this pipeline, we make the following novel contributions about the CWS vetting process and its effectiveness:
  1. 1.
    We find 17,021 potentially infringing extensions that are still published on the CWS. Our pipeline flags these extensions as extremely similar to known infringing extensions previously taken down from the store. Interestingly, 86% of these infringing items are republished extensions, i.e., extensions that are either identical or extremely similar to items that were previously taken down, suggesting that the CWS vetting process does not adequately learn from experience. We also find that 11% of the infringing extensions that remain published at the end of our crawl come from repeat offenders (publishers with known vetted items), suggesting that the CWS does not suspend these accounts according to their own policies. We also conduct a survival analysis for infringing extensions. Our results reveal that the process of identifying infringing extensions is excessively slow, taking months or even years from the publication of an infringing version to its removal from the store.
  2. 2.
    To validate our findings, we analyze extensions from the top infringing clusters with the most extensions and manually assigned them a type or kind of content. We find that 83% of the extensions in these clusters are New Tab Extensions (NTEs), which are items that override the webpage that loads when a new tab is opened. The cluster analysis also reveals the presence of spam and no content extensions, both vetted and still published in the CWS. In order to better understand the labeling process, we explore the relationship between the CWS vetting labels and the behavior of an extension. Our analysis finds that labels are somewhat inconsistent and unrelated to features or behaviors, especially in the case of NTEs.
  3. 3.
@@
    We scan the 5,647 extensions labeled as malware by the CWS in VirusTotal and compare the results. Remarkably, 95% of malware-labeled samples were not indexed by VirusTotal and 3% were known but have zero detections. This finding suggests that the majority of malicious extensions detected in the CWS are unknown to the threat intelligence community, and that for others there is a significant discrepancy between the criteria of CWS moderators and the detection capabilities of malware detection engines. Analysis of the malicious extensions for which we obtain detection reports reveals that the malware lineage analysis conducted by security vendors is very poor, with the family label missing in most cases and being very general in others.
@@
Disclosure. We reported to Google a subset of 180 infringing extensions found by our pipeline. All of them were manually reviewed to confirm the presence of a policy violation. Most importantly, this set includes 23 no content extensions published from corporate Google accounts, all of which were removed from the store shortly after our report. As of this writing, we have not received an official response.
@@
Tool and Research Artifacts. We release the source code of Fakeium, an instrumented JavaScript sandbox that can be used to extract API calls from browser extensions [42]. We also provide the list of extension pairs used as ground truth for evaluation, and metadata and high-dimensional embeddings for infringing extensions found by our pipeline at https://zenodo.org/records/10977708.
II Background
Figure 1: Data collection and analysis methodology pipeline.
@@ II-A Browser Extensions
Browser extensions are third-party programs that enhance or modify the functionality of the browser, with ad blockers being one of their most recognizable and widely-used examples. Extensions are written using a combination of HTML, CSS and JavaScript, and may contain other resources commonly found in web applications, such as images or fonts [39]. Extensions are packaged as CRX files, bundling all the source code and assets needed to run them into a single file, thus easing distribution. CRX packages are regular ZIP archives with an additional header that ensures the integrity and authenticity of its contents [49].
@@
As in the case of other popular platforms like Android, all extensions have a mandatory manifest.json file that dictates their functionality and behavior. This manifest is a JSON document that specifies basic metadata about the extension, like its name and version. Optionally, it may also declare permissions, service workers (which run code in the background), content scripts (which run code in open tabs), and overridden pages (e.g., to modify the default URL when opening a new tab), among other elements [10]. Two popular types of extensions are themes and New Tab Extensions (NTEs). A theme is a type of extension that changes the way the browser looks and does not contain any source code, i.e., all theme configuration is in the manifest itself. An NTE is an extension that changes the default webpage that will load every time a new tab is opened.
@@
Extensions communicate with the browser using Web APIs [40] and Extension APIs [21]. These APIs provide extensions with a broad range of powerful capabilities, allowing them to access and manipulate the Document Object Model (DOM) of webpages, and to access geolocation data, browsing history, and other sensitive information stored in the browser. Some of these APIs are restricted by default for security reasons, and can only be accessed by declaring the appropriate permissions in the manifest.
II-B CWS Program Policies
@@
CRX packages can be published on the Chrome Web Store (CWS), Google’s official distribution platform for browser extensions that run on Google Chrome [22]. Users of other Chromium-based browsers, such as Microsoft Edge, Brave, or Opera, can also install extensions from this source as they run the same compatible engine. In fact, Brave even recommends that users download extensions from this platform in the absence of their own official store [2]. As such, policing the content that is uploaded to the CWS is of great importance because of the impact that harmful extensions can have given the large audience of the store.
@@
The CWS requires all developers to comply with a set of policies that include security and privacy guidelines. These policies prohibit the publication of extensions with a malicious purpose and extensions that facilitate unauthorized access to restricted or copyrighted content [17]. They also ask developers not to obfuscate code or conceal functionality, to request the minimum set of permissions needed, and forbid submitting multiple extensions that provide duplicate or very little functionality [16, 15, 18]. According to Google’s policies, repeated violations of these policies will result in the suspension of the publisher account [19].
III Data Collection
@@
We implemented a purpose-specific crawler to download the resources shown in Figure 1, namely CRX packages and their associated metadata, from Google’s Chrome Web Store [22]. Our crawling strategy consists of three steps. First, we prepare an initial set of extensions from the store’s sitemap, the reels of popular items from each category, and the featured extensions from the home page. We then visit each extension page, extract metadata (e.g., name, description, number of installs, last modification date) and download its associated assets (i.e., CRX package, icon and tile). Finally, we look for non-visited extensions appearing as “related” or “recommended” in the store page of each visited extension and add them to the queue of items to crawl. Previous efforts demonstrate the effectiveness of this approach [9, 31], which takes inspiration from strategies successfully used to crawl the Google Play Store [65, 50].
@@
We run the crawler daily to detect new and removed extensions, and keep track of changes being pushed to the store such as publishers uploading an updated version of an existing extension. For simplicity, we set the visitor’s country code to the United States when crawling the store. Although developers can opt-out of publishing extensions in certain countries or even entire regions [14], this decision should not limit our coverage as we can still download any non-private CRX package regardless, and because the used sitemap contains publicly listed extensions worldwide. We only run our data collection pipeline once a day and rely on the HTTP caching headers provided by CWS to responsibly limit our impact on Google’s servers. We note that in October 2023, we transitioned to crawling the new version of the CWS, coinciding with its first public release [27]. We made this decision in preparation for the announced retirement of the previous version in January 2024. While we experienced no issues during this migration, we also kept crawling the legacy store for a few more weeks in the event something went wrong.
@@
In addition to crawling CRX packages and store metadata, we also fetch vetting labels issued by Google from a separate endpoint. These labels indicate whether an extension was taken down from the CWS instead of willingly being unpublished by its developer, as well as the reason behind the takedown (e.g., flagged as malware), if any. Vetting labels are used by Google Chrome since version 117 to protect users of vetted extensions by proactively disabling malware and displaying warnings in the browser extensions page [7]. As far as we can tell, this is the first study to crawl and analyze this data.
III-A Dataset
@@
We crawled the CWS for 4 years from March 2020 to March 2024. We compiled a historical dataset of 366,617 extensions, containing 902,532 different updates (or extension versions) that were available for download at some point in the store. Some of the extensions in our datasets were published as early as December 2009.111 While the CWS was officially launched in 2011, it existed prior to that date as evidenced by https://narkive.com/bEuDVJqN Of all these versions, we lack the CRX package for 28,405 paid extensions that did not change their pricing model after payments were deprecated in the CWS in mid-2020 [13]. Our dataset also contains 53,132 browser themes, which are distributed in the same form as extensions but do not have the ability to execute code. To the best of our knowledge, this is the most comprehensive dataset of browser extensions to-date, with over 156k more extensions than the previous large-scale work that crawled the CWS [48].
@@
Figure 2: Daily volume of published extensions (solid, left axis) and accumulated dataset size (dashed, right axis).
@@
Vetting Labels. Using the vetting labels that we crawl for removed extensions, we can infer whether they were unpublished by the developer or taken down by Google. Vetted extensions have a type of either Malware, Policy Violation, or Minor Policy Violation, whereas items that were unpublished at the request of their developers have a type of None.
Figure 3 shows the ratio of extension labels found in our dataset at a given point in time since we first started crawling. As of March 2024, 52% of all the extensions we collected have been removed from the store. More interestingly, 79% of the removed items are infringing extensions that were taken down. This preliminary characterization of the dataset motivates us to delve deeper into the dynamics of extension removals and the vetting process of the CWS.
Figure 3: Daily ratio of vetting labels provided by CWS.
IV SimExt: A Methodology for Finding Similar Extensions
While developing a fully automated vetting process is very challenging, having an automatic approach to detect similar extensions at scale is valuable to assist analysts, as it can be used to find items that resemble previously vetted ones or clones to currently published content. Our key working hypothesis is that if an extension is taken down from the store because it violates a program policy, then other extensions that are highly similar are candidates to be vetted as well. Even in those cases where the behavior of an extension is not necessary harmful, finding clones is still valuable given the CWS ban on repetitive content [15]. We acknowledge that legitimate extensions cloned by malicious actors are an exception to this rule. For example, the vetting of malicious extensions reusing MetaMask’s codebase should not result in MetaMask being taken down. We consider that these cases should be manually filtered out, and that a tool that provides detailed similarity reports between two extensions can be extremely helpful to assist the store vetting process. Section VI provides examples of published infringing extensions found by our tool to support this claim.
Since there is currently no tool that identifies similar extensions for a given input item, we designed and built SimExt, our own state-of-the-art solution for querying browser extensions based on similarity. We design it as a multi-step pipeline to automatically extract meaningful behavioral features from all extensions in our dataset and then group them into clusters based on their similarity. This approach is different from previous work that classifies extensions as either benign or malicious [28, 1, 48]. Moreover, to the best of our knowledge, our work is the first one to leverage Natural Language Processing (NLP) and vector embeddings to effectively and accurately find similar extensions at scale. Figure 1 provides a summary of the pipeline steps described below.
IV-A Feature Extraction